A silversmith mixed 25 g of a 60% silver alloy with 40 g of a 5% silver alloy. What is the percent concentration of the resulting alloy? (Round to the nearest tenth of a percent.)
%

Answer by mananth(16946)

percent ---------------- quantity
alloy 60 ---------------- 25 gms
alloy 5 ---------------- 40 gms
Total x ---------------- 65 gms
60*25+5* 40=65x
1500+200 =65x
1700=65x
/65
1700/65= x
26.15=x
26.5% mix